(y')2=(1-y^2)/(1-x^2) ; y=1/2 when x= 1
1
Expert's answer
2021-09-01T09:40:01-0400

"solve ~ (y^{\prime})^2= \frac{1-y^2}{1-x^2}, ~ y(1)=\frac{1}{2}, \\ 
y^{\prime} = \pm\sqrt{\frac{1-y^2}{1-x^2}}\\
\frac{dy}{\sqrt{1-y^2}}=\pm\frac{dx}{\sqrt{1-x^2}}\\
\int\frac{dy}{\sqrt{1-y^2}}=\pm\int\frac{dx}{\sqrt{1-x^2}}\\
\ln{|y+\sqrt{1-y^2}|}=\pm\ln{|x+\sqrt{1-x^2}|} +\ln|c| \\
\left[ 
 \begin{gathered} 
 y+\sqrt{1-y^2}=c(x+\sqrt{1-x^2}),\\
 y+\sqrt{1-y^2}=\frac{c}{(x+\sqrt{1-x^2})}.\\
 \end{gathered} 
\right.\\
y(1)=\frac{1}{2}.\\
\left[ 
 \begin{gathered} 
 \frac{1}{2}+\sqrt{1-(\frac{1}{2})^2}=c(1+\sqrt{1-1^2}),\\
 \frac{1}{2}+\sqrt{1-(\frac{1}{2})^2}=\frac{c}{(1+\sqrt{1-1^2})}.\\
 \end{gathered} 
\right.\\
\left[ 
 \begin{gathered} 
 \frac{1}{2}+\sqrt{\frac{3}{4}}=c,\\
 \frac{1}{2}+\sqrt{\frac{3}{4}}=c.\\
 \end{gathered} 
\right.\\
c=\frac{1+\sqrt{3}}{2}\\
answer: \\ \left[ 
 \begin{gathered} 
 y+\sqrt{1-y^2}=\frac{(1+\sqrt{3})}{2}(x+\sqrt{1-x^2}),\\
 y+\sqrt{1-y^2}=\frac{1+\sqrt{3}}{2(x+\sqrt{1-x^2})}.\\
 \end{gathered} 
\right.\\"